OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5).

OpenStudy (anonymous):

parallel lines have the same slope so m=3. You also have a given point (-10,2.5). Use point slope formula y-yo=m(x-xo) where (yo,xo) is your initial point. y-2.5=3(x+10) y=3x+32.5

OpenStudy (anonymous):

Oh, okay. That makes sense. Thank you!